OpenStudy (anonymous):

Please help with this word problem : On a clock, the minute hand is 4" long and the hour hand is 3" long. How fast is the distance b/t the tips os the hands changing at 9pm. answering in minutes/hour. Thank you!

OpenStudy (amistre64):

law of cosines might help ....

OpenStudy (amistre64):



OpenStudy (anonymous):

Im not sure how to proceed. ? I think it has to do with derivatives?

OpenStudy (dumbcow):

use law of cosines as @amistre64 said \[x^2 = 3^2 +4^2 - 2(3)(4) \cos \theta\] \[x^2 = 25 - 24 \cos \theta\] take derivative \[2x \frac{dx}{dt} = 24 \sin \theta \frac{d \theta}{dt}\] \[\frac{dx}{dt} = \frac{12 \sin \theta}{x} \frac{d \theta}{dt}\] Now we need to know how fast angle is changing each minute minute hand ---> moves 60 minutes around circle ---> 2pi = 60--> pi/30 radians/min hour hand ---> moves 12 hours or 720 minutes around circle --> 2pi=720 --> pi/360 The angle between them is the difference ---> pi/30 - pi/360 = 11pi/360 \[\theta(t) = \frac{11 \pi}{360} t\] \[\frac{d \theta}{dt} = \frac{11\pi}{360}\] given value of 9pm = 9 hrs = 540 min angle simplifies to pi/2 x = 5 plug values in \[\frac{dx}{dt} = \frac{12 \sin(\pi/2)}{5}*\frac{11\pi}{360}\] \[\frac{dx}{dt} = \frac{11\pi}{150}\]

OpenStudy (anonymous):

What units is that?

OpenStudy (anonymous):

I need to come up with inches per hour, and the number I came up with is a bit off from that.

OpenStudy (dumbcow):

oh i did inches per min .... just multiply by 60 to get in/hr

OpenStudy (anonymous):

11 pi /3?

OpenStudy (dumbcow):

\[\frac{11\pi}{150}*60 = \frac{22\pi}{5}\]

OpenStudy (anonymous):

Oh duh, ok, still not what I originally got but I see what I did wrong, thanks!

OpenStudy (dumbcow):

yw :)